# Multi-objective Bayesian Optimisation of Spinodoid Cellular Structures for Crush Energy Absorption

### v1

@@ 2.4 Effect of design parameters on performance indicators
Using the FEM simulations, performance indicators were extracted for a series of Design of Experiments (DoEs), wherein input design parameters were generated by quasi-random sampling of the Sobol’ Sequence [70]. This methodology ensures that any set of DoEs effectively cover the span of the available design space. Here, a mesh resolution of 30 and a wave number of $\lambda=15\pi$ have been chosen (See A.2, and A.3 for the choice of those specific values). The results of the one-at-a-time analysis have been presented in Fig. 6(a), which shows how individual parameters affect the performance indicators, assuming only one parameter is being varied at a time, with initial design being $\rho=0.3$ and $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$. An increase in relative density leads to a rise in both EA and PF, as expected due to the greater volume of solid material. However, the relationship between $\theta$ and the performance indices is complex. For instance, an increase in the performance indicators is observed for $\theta_{1}$ and $\theta_{2}$ until a value of 45 has been reached, with a drop in performance indicators for an increase in cone angles beyond 45. In contrast, an increase in performance indices is observed for $\theta_{3}$ values exceeding 60, whereas similar values of EA and PF are obtained for values below 60. When plotting all cone angles together, as shown in Fig. 6(b), $\theta_{1}$ and $\theta_{2}$ exhibit similar behaviour, whereas $\theta_{3}$ shows distinct variations. At a cone angle of 90, EA and PF converge to similar values, which can be attributed to isotropic structures generated at high cone angles.
Figure 6: (a) Visualisation of the one-at-a-time analysis, demonstrating how varying each individual design parameter affects the performance indicators. The initial design parameters were set at $\rho=0.3$ and $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$. Starting from the top-left and moving clockwise, the results show the effect of varying only the relative density, whilst keeping $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$, with $\rho$ values ranging from 0.3 to 0.6. For $\theta_{1}$, the relative density was held constant at $\rho=0.3$ and $\theta_{2}=\theta_{3}=15^{\circ}$. The same approach was applied for $\theta_{2}$ and $\theta_{3}$. Each subplot includes the topologies generated for specific design parameters. (b) The subplots display the impact on the performance indicators - EA on the left and PF on the right - when varying the different conical angles individually.
@@
In this study, two independent sensitivity analyses were conducted for each objective. Fig. 7 shows the results of applying Sobol’ sensitivity analysis to the spinodoid design space that is composed of four input parameters $(\rho,\theta_{1},\theta_{2},\theta_{3})$. Consequently, each pair of Sobol’ sequence points sampled necessitates the evaluation of six different points, as exemplified in Table 1. A total of 474 function evaluations were required to achieve converged sensitivity indices. The progression of these indices with the number of evaluations is depicted in A.4.1 with four design parameters, although further increasing the number of evaluations introduces numerical instabilities. The bar charts in Fig. 7 indicate a significant relationship between relative density and the performance indicators, with both EA and PF increasing as relative density increases. This is evidenced by the sensitivity indices S1, and ST, which are approximately one for relative density, while the indices for the remaining design parameters are close to zero. These results indicate that relative density is the most influential parameter on the performance indicators, while the other input parameters contribute very little. However, this does not imply that the different input parameters should be disregarded, as they play a critical role in influencing the anisotropy of structure, which is a key determinant of the performance of topologies under compression.
Figure 7: Bar charts illustrating the results of two independent Sobol’ sensitivity analyses for two performance indicators, EA and PF. These analyses were conducted using four inputs $(\rho,\theta_{1},\theta_{2},\theta_{3})$ with 474 model evaluations and the results are represented as sensitivity indices: S1 (First-order) and ST (Total-order), along with their confidence bounds.
@@
A total of 1,500 points were used to evaluate the Sobol’ sensitivity indices, increasing in increments of 5 points. The evolution of the sensitivity indices as the number of Sobol’ sequence samples increases is illustrated in A.4.2. Converged values of S1 and ST were obtained at approximately 1,200 evaluations, beyond which additional data had minimal effect on the indices. However, the sensitivity indices, along with their confidence bounds, are shown in Fig. 8 with a total of 1,500 evaluations. Unlike the previous sensitivity analysis, markedly different behaviour is observed. Focusing on the first-order sensitivity index, it is evident that all three inputs directly influence the performance indicators, with approximately 20% of the output variance being explained by the variance in a single input alone.
@@
Figure 8: Bar charts illustrating the results of two independent Sobol’ sensitivity analyses for two performance indicators, EA and PF. These analyses were conducted using three inputs $(\theta_{1},\theta_{2},\theta_{3})$ with 1,500 model evaluations, and the results are represented as sensitivity indices: S1 (First-order) and ST (Total-order), along with their confidence bounds.
@@ 3.1 Gaussian Processes
GPs can be trained using a given set of observations, $O=\left\{\textbf{x}_{i},\text{y}_{i}\right\}^{N}_{i=1}$, where $\textbf{x}_{i}\in\mathbb{R}^{q}$, given $N$ number of observations, and in a $q$ dimensional space. In addition, assuming noisy observations, $\text{y}_{i}=f(\textbf{x}_{i})+\epsilon_{i}\in\mathbb{R}$, where the noise, $\epsilon\sim\mathcal{N}(0,\bar{\sigma^{2}_{n}})$, is mean-centered and has a variance $\bar{\sigma^{2}_{n}}$. Consequently, the Gaussian distribution can then be obtained using
@@ 3.2 Scalarisation methods
Subsequently, both scalarisation techniques are used with a singular GP as a surrogate and an acquisition function such as the Expected Improvement (EI). In this study, the log-EI function has been utilised, which is an extension of the EI function, expressed as
@@ 4.1 BO setup
To demonstrate the implementation of the MOBO framework, two benchmark test problems were used: a) the Branin-Currin problem with added observation noise for each objective, and b) the ZDT1 with no additional observational noise. Both are maximisation problems. Five methods were tested, including the noiseless qEHVI function, to compare its effectiveness against its noisy counterpart qNEHVI. Additionally, a comparison was made between two different kernels - the RBF and Matern kernels - for calculating the covariance matrix. Before optimisation, six samples were generated from the design space to form the initial dataset, with the maximum number of iterations set at 200. The HV was used to assess the quality of the solutions provided by each method. The benchmark studies led to the following conclusions: 1) Regardless of the problem type, the Matern kernel outperformed the RBF kernel by capturing more intricate relationships between neighbouring points, allowing most methods to converge faster and achieve a total hypervolume closer to the maximum hypervolume. Consequently, the Matern kernel was selected for solving the spinodoid structure optimisation problem. 2) In noisy settings, the qNEHVI function outperformed qEHVI as expected. However, in the noiseless problem, both qEHVI and qNEHVI performed similarly, suggesting that qNEHVI is generally preferable. Thus, qNEHVI was one of the methods employed in the study of spinodoid cellular structures. Further details on the benchmark studies can be found in A.5.
@@ 4.2 Spinodoid problem setup
With the BO framework discussed and tested in Subsec. 4.1, it is crucial to emphasize the second part of the MOBO process, which involves integrating the MATLAB-ABAQUS interface to complete the optimisation cycle, as visualised in Fig. 10. The initial dataset for the BO process was created by sampling 50 data points using the Sobol’ sequence, which was subsequently evaluated. The pair plot in A.6 demonstrates the relationships between each design input and objective, highlighting a positive correlation between EA and PF.
@@ 4.2.3 MOBO workflow
In this study, sequential optimisation was performed with $q=1$. The stochastic nature of spinodoid topologies, due to GRFs and random initialisations, results in varying microstructures and randomness in the measured mechanical properties, introducing noise into the observations. This makes acquisition functions that can handle noisy data particularly suitable for this problem. However, to ensure repeatability, a seed can be set for the random number generator, in this case, it was set to rng(123) in MATLAB, producing consistent structures and eliminating noise. Nevertheless, in this study, the reference point was determined after multiple runs with various reference points, eventually settling on $\mathbf{r}=(-10,-0.3)$.
The optimisation loop was executed for a maximum of 250 iterations to ensure that a sufficiently populated Pareto front could be constructed, which also served as the stopping criterion. The full optimisation history can be found in A.7. The performance of the MOBO methods was evaluated against the NSGA-II algorithm by comparing the final hypervolume encompassed by the non-dominated solutions. The NSGA-II algorithm was implemented using the Python-based pymoo package, with the initial population derived from the initial dataset. The algorithm was run for 5 generations, resulting in a total of 250 evaluations ($50\times 5=250$) [103]. Simulated Binary Crossover (SBX) was used to generate offspring, while Polynomial Mutation was applied to introduce small variations in gene values, thus promoting diversity in subsequent generations. It should be noted that NSGA-II was used as a black box; further tuning of the crossover and mutation rates may improve the results. Specific hardware information, which was used to run the optimisation tasks on the cluster can be found in A.8.
@@ 5 Results
Fig. 14(a) illustrates the set of Pareto-optimal solutions derived from the qNEHVI method. The Pareto front includes 78 designs, each contributing uniquely to the overall hypervolume, Fig. 14(b) maps the rescaled Pareto-optimal points onto the original normalised objective plane, also displaying the initial dataset for comparison. Compared to the initial dataset, the MOBO method has identified several optimal designs that exhibit reduced PF while preserving the EA capability. Moreover, qNEHVI has uncovered previously undiscovered optimal points at the higher end of the EA range, particularly in the upper-right region of the normalised PF-EA graph. Six optimal points were selected to demonstrate the robustness of these designs, with Fig. 14(c) showcasing the generated topologies, featuring various microstructures with differing relative densities. The F-D behaviour of each selected design under compression is plotted in Fig. 14(d), revealing distinct performance characteristics. It is important to note that high PF values have not been observed in the optimal solutions, which are predominantly found in anisotropic structures, such as the ‘Columnar’-type demonstrated in the experimental validation shown in Fig. 4. Moreover, the optimal designs predominantly feature anisotropic structures, particularly those oriented perpendicular to the crushing direction such as ‘Lamellar’-type, as well as structures with mixed geometries that lack distinct characteristics or any correlation with conical angles. These are examples of bending-dominated structures, where loads are primarily resisted through bending and flexural mechanisms. While these structures tend to have lower stiffness and strength due to localised deformations that can lead to early failure through buckling or yielding, the bending mechanisms enable stress distribution across the structure. This allows the load to be spread over a larger area as the structure deforms more easily, effectively delaying catastrophic failure, which results in such structures having higher EA. It should also be noted that these solutions were deemed to not have undergone densification, which is also evident in the F-D curves of the selected cases.

### v2

@@ 2.4 Effect of design parameters on performance indicators
Using the FEM simulations, performance indicators were extracted for a series of Design of Experiments (DoEs), wherein input design parameters were generated by quasi-random sampling of the Sobol’ Sequence [70]. This methodology ensures that any set of DoEs effectively cover the span of the available design space. Here, a mesh resolution of 30 and a wave number of $\lambda=15\pi$ have been chosen (See A.3, and A.5 for the choice of those specific values). The results of the one-at-a-time analysis have been presented in Fig. 6(a), which shows how individual parameters affect the performance indicators, assuming only one parameter is being varied at a time, with initial design being $\rho=0.3$ and $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$. An increase in relative density leads to a rise in both EA and PF, as expected due to the greater volume of solid material. However, the relationship between $\theta$ and the performance indices is complex. For instance, an increase in the performance indicators is observed for $\theta_{1}$ and $\theta_{2}$ until a value of 45 has been reached, with a drop in performance indicators for an increase in cone angles beyond 45. In contrast, an increase in performance indices is observed for $\theta_{3}$ values exceeding 60, whereas similar values of EA and PF are obtained for values below 60. When plotting all cone angles together, as shown in Fig. 6(b), $\theta_{1}$ and $\theta_{2}$ exhibit similar behaviour, whereas $\theta_{3}$ shows distinct variations. At a cone angle of 90, EA and PF converge to similar values, which can be attributed to isotropic structures generated at high cone angles.
Figure 6: (a) Visualisation of the one-at-a-time analysis, demonstrating how varying each individual design parameter affects the performance indicators. The initial design parameters were set at $\rho=0.3$ and $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$. Starting from the top-left and moving clockwise, the results show the effect of varying only the relative density, whilst keeping $\theta_{1}=\theta_{2}=\theta_{3}=15^{\circ}$, with $\rho$ values ranging from 0.3 to 0.6. For $\theta_{1}$, the relative density was held constant at $\rho=0.3$ and $\theta_{2}=\theta_{3}=15^{\circ}$. The same approach was applied for $\theta_{2}$ and $\theta_{3}$. Each subplot includes the topologies generated for specific design parameters. (b) The subplots display the impact on the performance indicators - EA on the left and PF on the right - when varying the different conical angles individually. It should be noted that these structures were also crushed along the $x_{3}$ direction.
@@
In this study, two independent sensitivity analyses were conducted for each objective. Fig. 7 shows the results of applying Sobol’ sensitivity analysis to the spinodoid design space that is composed of four input parameters $(\rho,\theta_{1},\theta_{2},\theta_{3})$. Consequently, each pair of Sobol’ sequence points sampled necessitates the evaluation of six different points, as exemplified in Table 1. A total of 474 function evaluations were required to achieve converged sensitivity indices. The progression of these indices with the number of evaluations is depicted in A.7.1 with four design parameters, although further increasing the number of evaluations introduces numerical instabilities. The bar charts in Fig. 7 indicate a significant relationship between relative density and the performance indicators, with both EA and PF increasing as relative density increases. This is evidenced by the sensitivity indices S1, and ST, which are approximately one for relative density, while the indices for the remaining design parameters are close to zero. These results indicate that relative density is the most influential parameter on the performance indicators, while the other input parameters contribute very little. However, this does not imply that the different input parameters should be disregarded, as they play a critical role in influencing the anisotropy of structure, which is a key determinant of the performance of topologies under compression.
Figure 7: Bar charts illustrating the results of two independent Sobol’ sensitivity analyses for two performance indicators, EA and PF. These analyses were conducted using four inputs $(\rho,\theta_{1},\theta_{2},\theta_{3})$ with 474 model evaluations and the results are represented as sensitivity indices: S1 (First-order) and ST (Total-order), along with their confidence bounds. The results reveal that $\rho$ exhibits extremely high sensitivity, while the $\theta$ inputs show nearly negligible sensitivity.
@@
A total of 1,500 points were used to evaluate the Sobol’ sensitivity indices, increasing in increments of 5 points. The evolution of the sensitivity indices as the number of Sobol’ sequence samples increases is illustrated in A.7.2. Converged values of S1 and ST were obtained at approximately 1,200 evaluations, beyond which additional data had minimal effect on the indices. However, the sensitivity indices, along with their confidence bounds, are shown in Fig. 8 with a total of 1,500 evaluations. Unlike the previous sensitivity analysis, markedly different behaviour is observed. Focusing on the first-order sensitivity index, it is evident that all three inputs directly influence the performance indicators, with approximately 20% of the output variance being explained by the variance in a single input alone.
@@
Figure 8: Bar charts depicting the outcomes of two separate Sobol’ sensitivity analyses for the performance indicators EA and PF. These analyses were performed using three inputs $(\theta_{1},\theta_{2},\theta_{3})$ across 1,500 model evaluations while keeping the $\rho$ constant at $\rho=0.3$. The results, presented as sensitivity indices—S1 and ST—include corresponding confidence bounds, showing an even distribution of sensitivities among the inputs $\theta$.
@@ 3.1 Gaussian Processes
GPs can be trained using a given set of observations, $O=\left\{\textbf{x}_{i},\text{y}_{i}\right\}^{N}_{i=1}$, where $\textbf{x}_{i}\in\mathbb{R}^{q}$, given $N$ number of observations, and in a $q$ dimensional space. In addition, assuming noisy observations, $\text{y}_{i}=f(\textbf{x}_{i})+\epsilon_{i}\in\mathbb{R}$, where the noise, $\epsilon\sim\mathcal{N}(0,\bar{\sigma^{2}_{n}})$, is modelled as mean-zero Gaussian noise with fixed variance $\bar{\sigma^{2}_{n}}$. This approach simplifies the noise modelling by treating the results from the FEM simulations as the ‘ground truth’. However, in practice, the noise term may not have a mean of zero ($\mu\neq 0$), as it encompasses two types of uncertainty: aleatoric (data) uncertainty and epistemic (model) uncertainty. Aleatoric uncertainty arises from random variations in observations, such as sensor noise or experimental imperfections, and is inherently irreducible. Epistemic uncertainty, on the other hand, stems from limitations or approximations in the computational models or objective function calculations, which may not fully capture real-world conditions due to idealised assumptions or incomplete knowledge of the system. As a result, this work focuses on aleatoric uncertainty and models the noise as mean-zero Gaussian noise ($\epsilon\sim\mathcal{N}(0,\bar{\sigma^{2}_{n}})$) for simplicity. To better address epistemic uncertainty, the noise model could be refined using prior knowledge, such as FEM validation results (e.g., Fig. 4), to determine the noise parameters (mean and variance) more accurately.
Consequently, the Gaussian distribution can then be obtained using
@@ 3.2 Scalarisation methods
Subsequently, both scalarisation techniques are used with a singular GP as a surrogate and an acquisition function such as the Expected Improvement (EI). EI is a widely used acquisition function and one of the earliest methods developed for efficient global optimisation tasks [98, 99]. However, in this work, an extended version of the EI function is employed to address the issue of vanishing gradients that can arise when optimising improvement-based acquisition functions [100]. It is important to note that the choice of acquisition function is largely heuristic, serving as a strategy to navigate the design space and map it onto a belief space. This choice offers a practical and efficient approach to navigating the design space, particularly in cases like this, where the influence of design variables on the objective functions (energy absorption and peak force) varies significantly. In this study, the log-EI function has been utilised, which is an extension of the EI function, expressed as
@@ 4.1 BO setup
To demonstrate the implementation of the MOBO framework, two benchmark test problems were used: a) the Branin-Currin problem with added observation noise for each objective, and b) the ZDT1 with no additional observational noise. Both are maximisation problems. Five methods were tested, including the noiseless qEHVI function, to compare its effectiveness against its noisy counterpart qNEHVI. Additionally, a comparison was made between two different kernels - the RBF and Matern kernels - for calculating the covariance matrix. Before optimisation, six samples were generated from the design space to form the initial dataset, with the maximum number of iterations set at 200. The HV was used to assess the quality of the solutions provided by each method. The benchmark studies led to the following conclusions: 1) Regardless of the problem type, the Matern kernel outperformed the RBF kernel by capturing more intricate relationships between neighbouring points, allowing most methods to converge faster and achieve a total hypervolume closer to the maximum hypervolume. Consequently, the Matern kernel was selected for solving the spinodoid structure optimisation problem. 2) In noisy settings, the qNEHVI function outperformed qEHVI as expected. However, in the noiseless problem, both qEHVI and qNEHVI performed similarly, suggesting that qNEHVI is generally preferable. Thus, qNEHVI was one of the methods employed in the study of spinodoid cellular structures. Further details on the benchmark studies can be found in A.8.
@@ 4.2 Spinodoid problem setup
With the BO framework discussed and tested in Subsec. 4.1, it is crucial to emphasise the second part of the MOBO process, which involves integrating the MATLAB-ABAQUS interface to complete the optimisation cycle, as visualised in Fig. 10. The initial dataset for the BO process was created by sampling 50 data points using the Sobol’ sequence, which was subsequently evaluated. The pair plot in A.9 demonstrates the relationships between each design input and objective, highlighting a positive correlation between EA and PF.
@@ 4.2.3 MOBO workflow
In this study, sequential optimisation was performed with $q=1$. The stochastic nature of spinodoid topologies, due to GRFs and random initialisations, results in varying microstructures and randomness in the measured mechanical properties, introducing noise into the observations. This makes acquisition functions that can handle noisy data particularly suitable for this problem. However, to ensure repeatability, a seed can be set for the random number generator, in this case, it was set to rng(123) in MATLAB, producing consistent structures and eliminating noise. Nevertheless, in this study, the reference point was determined after multiple runs with various reference points, eventually settling on $\mathbf{r}=(-10,-3)$.
The optimisation loop was executed for a maximum of 250 iterations to ensure that a sufficiently populated Pareto front could be constructed, which also served as the stopping criterion. The full optimisation history can be found in A.10. The performance of the MOBO methods was evaluated against the NSGA-II algorithm by comparing the final hypervolume encompassed by the non-dominated solutions. The NSGA-II algorithm was implemented using the Python-based pymoo package, with the initial population derived from the initial dataset. The algorithm was run for 5 generations, resulting in a total of 250 evaluations ($50\times 5=250$) [106]. Simulated Binary Crossover (SBX) was used to generate offspring, while Polynomial Mutation was applied to introduce small variations in gene values, thus promoting diversity in subsequent generations. It should be noted that NSGA-II was used as a black box; further tuning of the crossover and mutation rates may improve the results. Specific hardware information, which was used to run the optimisation tasks on the cluster can be found in A.11.
@@ 5 Results
Fig. 14(a) illustrates the set of Pareto-optimal solutions derived from the qNEHVI method. The Pareto front includes 78 designs, each contributing uniquely to the overall hypervolume, Fig. 14(b) maps the rescaled Pareto-optimal points onto the original normalised objective plane, also displaying the initial dataset for comparison. Compared to the initial dataset, the MOBO method has identified several optimal designs that exhibit reduced PF while preserving the EA capability. Moreover, qNEHVI has uncovered previously undiscovered optimal points at the higher end of the EA range, particularly in the upper-right region of the normalised PF-EA graph. Six optimal points were selected to demonstrate the robustness of these designs, with Fig. 14(c) showcasing the generated topologies, featuring various microstructures with differing relative densities. The F-D behaviour of each selected design under compression is plotted in Fig. 14(d), revealing distinct performance characteristics.
For instance, the reduction in normalised PF for the optimal points, such as the purple diamond (optimal point 4), corresponds to a decrease from 0.19 to 0.17, which translates to an approximate reduction in peak force of 1,610 N when normalised by the peak force of 80,476 N from compressing a cube of size $40\times 40\times 40\ \text{mm}^{3}$. Even small reductions in normalised PF can have meaningful implications, particularly in applications where performance or safety is sensitive to force thresholds. For example, in structural or collision scenarios, lower PF values can reduce material stress, enhance energy dissipation, and improve durability or safety. While the magnitude of the reduction may seem modest, such improvements often contribute to achieving more optimal designs or meeting stringent performance criteria. This highlights the value of even incremental advancements achieved through optimisation.
The absence of high PF values in the optimal solutions highlights the effectiveness of the optimisation process in balancing competing objectives. The optimal designs predominantly feature geometries with mixed values of $\theta$, which lack distinct characteristics or a clear correlation with conical angles. This outcome reflects the inherent flexibility of the MOBO approach, which identifies a diverse set of Pareto-optimal solutions by varying design variables to achieve desired trade-offs between EA and PF. As a result, the optimisation process systematically navigates the Pareto frontier, favouring designs that maximise energy absorption while keeping peak forces within acceptable limits.
These are examples of bending-dominated structures, where loads are primarily resisted through bending and flexural mechanisms. While these structures tend to have lower stiffness and strength due to localised deformations that can lead to early failure through buckling or yielding, the bending mechanisms enable stress distribution across the structure. This allows the load to be spread over a larger area as the structure deforms more easily, effectively delaying catastrophic failure, which results in such structures having higher EA. It should also be noted that these solutions were deemed to not have undergone densification, which is also evident in the F-D curves of the selected cases.
@@
6 Optimisation with three variables
The primary approach to obtaining Pareto-optimal solutions in the spinodoid design space involves optimising all four design variables simultaneously. This ensures globally optimal results without introducing any prior assumptions, allowing for a comprehensive exploration of the design space. Such an approach also captures complex interactions between variables, leading to a more diverse and well-populated Pareto front.
However, scenarios exist where efficiency gains can be achieved by leveraging insights from the sensitivity analysis highlighted in Subsec. 2.4. Specifically, the highly sensitive parameter, $\rho$, can be fixed prior to the optimisation based on performance criteria. For example, to obtain a set of Pareto-optimal solutions targetting the lowest weight, $\rho=0.3$ could be set before generating the initial dataset. This strategy reduces computational cost in two ways. First, this strategy reduces the dimensionality of the problem, allowing the optimisation process to focus on the remaining, less sensitive variables, $\theta$’s allowing convergence to occur faster. Second, the main contributor to the increased optimisation time is incurred when evaluating the next point recommended by the acquisition function. As the computation time for simulations scales with the relative density, a low value for $\rho$ would speed up the computation time to complete a single iteration i.e. time taken to conduct FE analysis on a single point, thus significantly reducing the time to complete the optimisation task. However, the opposite is also true, if a large value of $\rho$ is chosen, the computation time will significantly increase due to a greater number of elements required to mesh the geometry as well as the difference in mechanical behaviour.
Following the successful implementation of the qNEHVI method with four variables, a three-variable optimisation was conducted with the relative density fixed at $\rho=0.3$ prior to the optimisation. The Sobol’ sequence was again used to generate an initial dataset of 50 points, and the optimisation process was run for 250 iterations. Detailed optimisation history is provided in A.12. The evolution of the hypervolume difference across the 250 iterations is shown in Fig. 16(a). This figure compares the hypervolume of the three-variable optimisation ($\rho=0.3$ fixed) with that of the four-variable optimisation, isolating all Pareto-optimal solutions with $\rho=0.3$ i.e. $\text{HV}(\mathcal{P}_{\text{3var}})-\text{HV}(\mathcal{P}_{\text{4var}}[\rho=0.3])$, where $\text{HV}(\mathcal{P}_{\text{3var}})$ is the hypervolume of Pareto-optimal solutions obtained from three-variable optimisation, with $\rho=0.3$ fixed prior to the optimisation and $\text{HV}(\mathcal{P}_{\text{4var}}[\rho=0.3])$ is the hypervolume of Pareto-optimal solutions obtained from the four-variable optimisation, but filtered a posteriori to include only those solutions satisfying $\rho=0.3$ (Illustrated in Fig. 15(a)(i)). The positive difference observed in the initial iterations indicates that the three-variable optimisation outperforms the conditional solutions from the four-variable optimisation from the very first iteration. By focusing on the less sensitive parameters, the three-variable approach exploits correlations between the $\theta$’s and the objectives more effectively, leading to better hypervolume performance from early on. Furthermore, the three-variable method continues to demonstrate improvements as the optimisation progresses, highlighting its ability to efficiently refine the design space.
The scaled Pareto-optimal solutions obtained from the three-variable optimisation are shown in Fig. 16(b), with 26 points forming the Pareto-front. Similar to the four-variable optimisation, six optimal points on the Pareto front are marked with different coloured symbols. Unlike the Pareto front in Fig. 14(a), the three-variable solutions occupy a smaller portion of the objective space due to the lower relative density, with $\overline{\mathrm{PF^{\prime}}}$ ranging from -0.08 to approximately 0, and $\overline{\mathrm{EA^{\prime}}}$ ranging from -9 to -2. Despite the reduced objective space, the three-variable optimisation discovers new solutions by varying the $\theta$’s, achieving improved EA values while maintaining similar PF levels, which four-variable optimisation found by varying the $\rho$ that improves the EA. In some cases, however, the four-variable optimisation yields solutions that outperform those from the three-variable optimisation.
Fig. 16(c) presents the unscaled Pareto-optimal points in the normalised $\overline{\mathrm{PF}}$-$\overline{\mathrm{EA}}$ plane alongside the initial dataset. The initial Sobol’ sampling caused the evaluated objectives to cluster in the top-right region of the objective space, making it densely populated and reducing uncertainty in the Gaussian process regression (GPR) model in that region. In contrast, the sparsely explored bottom-left region exhibited higher uncertainty, which the acquisition function identified as an opportunity to improve the hypervolume. This led to a focused exploration in the bottom-left region of the $\overline{\mathrm{PF}}$-$\overline{\mathrm{EA}}$ plane, while the optimisation process effectively leveraged the initially sampled points in the top-right, incorporating them into the Pareto-optimal solutions, as indicated by the overlapping initial data and Pareto optimal solution markers. Illustration of topologies that result in objective values shown by the coloured markers have been shown in Fig. 16(d). Similar to the optimal design parameters obtained from four-variable optimisation, the structures generated here are also bending-dominated. These optimal solutions feature geometries with mixed values of $\theta$, leading to configurations that inherently facilitate rotational movement or the generation of bending moments under axial loading. This behaviour contrasts sharply with stretching-dominated structures, such as the ‘Columnar’ topology, which rely on direct axial force transfer. The bending-dominated designs resemble isotropic, bone-like structures, characterised by interconnected, irregular patterns that distribute loads over multiple pathways, enabling higher flexibility and energy absorption. Such geometries reflect nature-inspired design principles, where complex structural arrangements optimise mechanical performance by balancing strength, stiffness, and ductility. In conclusion, by predefining the relative density, the optimisation process more effectively maps the structure-property relationship between the cone angles and the objectives. This contrasts with the four-variable optimisation, where improvements in the Pareto front were primarily driven by initially adjusting the density to increase or decrease the objectives. The predefined density approach enables a more detailed and precise mapping between $\theta$’s and the objectives, as well as reducing the overall computation time from an average of 10-14 days as reported for the four-variable case down to approximately two days to complete optimisation up to 250 iterations, primarily driven by the lower cost of evaluating each recommended sample point.
Figure 16: Pareto-optimal solutions obtained from optimisation with three variables, $\theta$’s, conducted with a fixed relative density of $\rho=0.3$ using the qNEHVI method. Subplot (a) represents the difference in hypervolume between the Pareto-optimal solutions filtered post-optimisation to include only those with $\rho=0.3$, derived from a four-variable optimisation (as shown in Fig. 15(a)(i)), and the Pareto-optimal solutions obtained from a three-variable optimisation. The hypervolume difference is quantified as $\text{HV}(\mathcal{P}_{\text{3var}})-\text{HV}(\mathcal{P}_{\text{4var}}[\rho=0.3])$. This comparison is evaluated over 250 iterations, with the solid central line representing the median and the thinner lines at the top and bottom indicating the 25th and 75th percentiles, respectively. Subplot (b) presents Pareto-optimal points scaled using Eq. (27) obtained from three-variable and four-variable optimisation, with example geometries highlighted by markers of different colours shown in the insert. Subplot (c) displays the unscaled, normalised objectives, where black cross markers denote Pareto-optimal points, and red circular markers highlight the initial dataset. Subplot (d) visualises the topologies corresponding to specific objective values, as indicated by their respective markers. The values shown beneath each image represent the $\theta$’s required to generate the structures at $\rho=0.3$.
It should also be noted that due to the heavy influence of $\rho$ in four-variable optimisation, the acquisition function tends to prioritise $\rho$ and recommend values at the extremes of its design space, while the $\theta$’s are sampled in a more stochastic manner without specifically targetting the extreme ends of their ranges. This behaviour reflects the sensitivity disparity between $\rho$ and $\theta$’s, where $\rho$ drives most of the changes in the objectives, leading to less focused exploration of the $\theta$’s. To address this imbalance, standardisation was applied to approximately normalise the variables based on their order of magnitude, ensuring that they are of comparable scale. While the exact bounds of the variables are not fully known in the context of optimisation, this scaling helps mitigate the disparity in sensitivity and encourages more balanced exploration across all variables. The evolution of the recommended sample points over the course of the three-variable and four-variable optimisation can be found in A.13. Consequently, the four-variable optimisation process may require additional iterations to thoroughly explore and optimise the less sensitive variables, further contributing to the increased computational cost.
7 Discussion